OpenStudy (anonymous):

Given vector v with initial point: (-6,4,-2) and terminal point: (1,-1,3) Find a unit vector in the direction of v.

OpenStudy (amistre64):

last - start gives you directionality

OpenStudy (amistre64):

divide by length to get it to unit length

OpenStudy (amistre64):

3ft/3 = 1 12 inches/12 = 1 58 miles /58 = 1 \[\left|\frac{\vec v}{|v|}\right|=1\]

OpenStudy (anonymous):

My teacher gave the answer sqrt(11)/33 (7,-5,5) how would I get that (this was on a test I failed, I just want to know how to do it right)

OpenStudy (anonymous):

<1-(-6),-1-4,3-(-2)>=<7,-5,5> which is the vector. To find it's magnitude you do \[\sqrt{7^{2}+-5^{2}+5^{2}}\]

OpenStudy (anonymous):

a unit vector is the (vector/mag of the vector)

OpenStudy (anonymous):

That makes sense. I didn't know what to find the magnitude of

OpenStudy (anonymous):

Although the mag is 3(11^1/2)

OpenStudy (anonymous):

Yeah that's what I got too. Not sure how to get (11^1/2)/33

OpenStudy (anonymous):

I would mention that to him so you can get your points back on the exam!

OpenStudy (anonymous):

I'll ask tomorrow, but wont get any points back since I don't even know why I tried doing, but it was nowhere near the right answer. thanks for the help though :)